OpenStudy (anonymous):

Find the slope-intercept equation for the line with the indicated slope and y-intercept... slope 5/7; y-intercept 0,4

OpenStudy (aroub):

y=5/7x+4

OpenStudy (anonymous):

how did you come to that answer

OpenStudy (aroub):

y=mx+b where m is the slope and b is the y-intercept so u just substitute the numbers u have to y=mx+b u have 5/7 for m and 4 for the y-intercept so it becomes y=5/7x+4

OpenStudy (anonymous):

let us use the general equation of a line y=mx+c from the question m=slope=5/7 c=y-intercept=4 therefore the equation is y=5/7x+4 7y=5x+28 5x-7y+28=0 this is the equation of the line

OpenStudy (anonymous):

Thank you